Remember if the question says “what is the equation for the area of a circle?” Enter pi*r^2 not A=pi*r^2.

For frequency write nu.

For wavelength write lambda. Use RH for the Rydberg constant expressed in energy units (not wavelength), ie., RH = 2.18e-18J. Use h for Planck’s constant andc for the speed of light.

Note that 1 divided by b*c should be written 1/(b*c)


B. What is the equation for the energy of 1 photon when the calculation must be done using the frequency of light?


C. If the frequency of light is known what equation gives the wavelength of light?


D. What equation gives the wavelength of a particle?


F. What equation gives the energy of the n=5 level of the hydrogen atom?


G. What is the equation for the frequency of light when the Hydrogen atom makes a transition from n=7 to n=2? Evaluate everything but RH and h


1
Expert's answer
2021-02-12T05:02:14-0500

B). E = h*v)

C). c = v*lambda

v = c/lambda

D). lambda = c/v

F). Energy = RH/(5)² = (2.18×10-18)/25

= 8.72×10-20

G).( 1/lambda )= 109677( "\frac{1}{n_1^2}" - "\frac{1}{n_2^2}" )

n1 = 2 , n2 = 7

( 1/lambda ) = 109677 ( "\frac{1}{4}" - "\frac{1}{49}" )

( 1/lambda ) = 109677 × 0.229

( 1/lambda ) = 25116.03

lambda = 1/25116.03

lambda = 0.0000398 = 3.98 × 10-5